+3q +4q 0 Two pucks are placed on a frictionless table as shown above. Each puck has mass m. A positive charge of +3go is placed on the left charge and a charge of +4go is placed on the right puck. The pucks are released at rest and the distance D between the two pucks. ust after they are released? just after they are released? acceleration of the right pucks of 2.000 Gees (a) For each puck, what is the magnitude and direction of the electric force at the instant (b) For each puck, what is the magnitude and direction of the acceleration at the instant (c) If m = 1.300 kg and D-0.512 meters, what is the value of go that would yield an
Answer #1

Apply coulomb's law of electric force between two charges to solve the problem

formula

F = k q1 q2/r^2

------------------------------------------------------------

(a)

F = k ( 3q0)( 4q0)/D^2

= 12 k qo^2/D^2

toward right side

(b)

Apply Newton second law

F = ma

ma = k ( 3q0)( 4q0)/D^2

a = 12 k qo^2/D^2 * m

(c)

qo = sqrt aD^2 m/12 k

= sqrt (2(9.8)(0.512)^2(1.3)/12 (9* 10^9)

=7.86 * 10^-6 C